OpenStudy (anonymous):

solve for x. write both solutions, separated by a comma. 2x^2-7x+5=0

OpenStudy (anonymous):

quadratic formula, -b+/-sqrt(4ac-b^2)/(2a), standard equation, Ax^2+Bx+C, A=2,B=7, C=5, substitute values into quadratic formula and execute.

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

HELPPP???

OpenStudy (anonymous):

Did you take a look at my solution?

OpenStudy (anonymous):

\[ 2x^2 - 7x + 5 =0\\\\ \text{ Break the Middle term into two numbers , A and B, } \\ \text{ that, when added, equal -7 and when multiplied equal 5 } \\\\ 2x^2 - 2x -5x + 5\\\\ 2x( x -1) - 5x ( x -1 ) \\ ( 2x -5 ) ( x - 1) \\ x = \frac{ 5}{ 2 }\ \ \ and \ \ \ x = 1\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

uhhh, welcome